OpenStudy (logan13):

Lila earns $8.75/h at her job. How many hours must Lila work to earn $105? Complete the equation to model the problem. Use h to represent the unknown value.

OpenStudy (logan13):

8.75=

OpenStudy (anonymous):

12 i think

OpenStudy (logan13):

=8.75 sorry

OpenStudy (logan13):

its 2 numbers

OpenStudy (anonymous):

it should be 12 hrs to get to 105

OpenStudy (anonymous):

8.75h = 105 The above is the equation. \( \frac{8.75h}{8.75}= \frac{105}{8.75} \) \( h= \frac{105}{8.75} \) \( h= 12 \)